OpenStudy (anonymous):

I am really looking for someone to guide me through this. If you just want to give me the answer that's fine, but I would much rather know how to do this one. The equation for a toy car is given by x(t)=.5t-3.5. Find the position at 4 seconds. At what instant is the position zero for x(t) =.5t-3.5 and for x-4. Also, how long will it take for the position to reach 12 meters.

OpenStudy (anonymous):

x(t) is the distance x of time x(t) = .5t -3.5. therefore, at time t = 4 seconds x(4) = .5(4) - 3.5 = -1.5 meters at what point is the car at position 0 or x(t) = 0 = .5t - 3.5 .5t = 3.5 t = 7 seconds

OpenStudy (anonymous):

Same equations. That makes no sense. I was sure this LAB wants me to use different equations. Maybe its tricking me.

OpenStudy (anonymous):

or maybe I don't know what I am talking about, thanks man..lol

OpenStudy (anonymous):

nm I know what I am doing wrong!

OpenStudy (anonymous):

verify this for me if you look at this again! y=mx+b therefore m=velocity and b= position. therefore, original equation is x(t)=.5(t)-3.5. If I am creating an equation for 4 seconds, then the 3.5 would become 4 in the original equation correct?

OpenStudy (anonymous):

so the next equation would be x(t)-.5(t)-4. Then evaluate zero at t=4?

OpenStudy (anonymous):

err = not -